Calculate the quantity of oxygen in a bucket of fresh water (volume10litre) at two different altitudes .sea level and 5000 m.a.s.l use Henry's law and assume a temperature of 15°c



1
Expert's answer
2020-08-17T08:58:00-0400

Henry's law:


"x = \dfrac{p}{H}"

where "x" is the mole fraction of the oxigen in the water, "p" is the partial oxygen preassure in air (in atm) and "H = 3.64\times 10^{4}" is Henry's constant for oxygen at 15°c.

1. The partial oxygen preassure in air at sea level is:


"p_0 = 1\space atm\times 0.21 = 0.21\space atm"

where "0.21" represents the fraction of oxygen in air.

Thus:


"x_0 = \dfrac{0.21}{3.64\times 10^{4}} \approx 5.8\times 10^{-6}"

One liter of water contains 55.56 moles of water. Thus, 10 liters will contain 555.6 moles. Thus, the quantity of oxygen will be:


"n_0 = x_0\times 555.6 \space moles = 5.8\times 10^{-6}\times 555.6\space moles \approx 3.22\times 10^{-3}\space moles"

2. The partial oxygen preassure in air at 5000 m.a.s is:


"p_0 = 0.53\space atm\times 0.21 = 0.11\space atm"

where "0.21" represents the fraction of oxygen in air.

Thus:


"x_0 = \dfrac{0.11}{3.64\times 10^{4}} \approx 3.02\times 10^{-6}"

One liter of water contains 55.56 moles of water. Thus, 10 liters will contain 555.6 moles. Thus, the quantity of oxygen will be:


"n_0 = x_0\times 555.6 \space moles = 3.02\times 10^{-6}\times 555.6\space moles \approx 1.68\times 10^{-3}\space moles"

Answer. Sea level: 3.22*10^-3 moles, 5000 m.a.s: 1.68*10^-3 moles.